A) Suppose A is a 3 x 3 matrix such that det (A) = 1/125 . Find det (5A^-1)


B)Suppose A is a 3 x 3 matrix such that det (A) = 5. Determine det [A^4]
1
Expert's answer
2020-05-13T19:43:15-0400

A) "\det(5A^{-1})=5^3\det(A^{-1})=\frac{5^3}{\det A}=5^6"


B) "\det(A^4)=(\det A)^4=5^4"